# Ultracoherent superconducting cavity-based multiqudit platform with error-resilient control

### v2

@@ V Conclusion
The elliptical-shaped 2-cell cavity architecture demonstrated in this work can be naturally scaled by increasing the number of cells to realize larger multimode systems. These can be configured into cascaded random access quantum memories [77] with order-of-magnitude improvements in coherence times, where each two-cell module serves as a cache memory. Improved surface treatments [17, 78] can further improve quality factors by an order of magnitude. Additionally, these two-cell devices can function as modular building blocks for a scalable quantum module, leveraging quantum-state router [79] and state-of-the-art tunable coupling devices [7, 6, 8] for the exchange of quantum information and the distribution of entanglement. This setup can be further scaled by employing remote interconnection between the scalable modules, and even across different cryogenic systems, facilitating the development of a distributed quantum network with improved coherence and transformative computational power.
@@ Appendix A System Hamiltonian parameters
Values of the system parameters are shown in Table A1. The inverse-Purcell limit on the cavity mode lifetime is determined by both the relaxation as well as the dephasing channel of the transmon, which can be approximated as [15]
@@
From this, we estimate decay times of $T^{\mathrm{Purcell}}_{1a}=1.17\,$s and $T^{\mathrm{Purcell}}_{1b}=0.34\,$s, far exceeding the bare Alice and Bob lifetime limited by the cavity’s intrinsic material loss. It’s worth noting that this approximation assumes the same dephasing rate measured from the Hahn-echo experiment for the actual dressed-dephasing-induced cavity decay process. In reality, such a decay process is sensitive to dephasing noises at the detuning between the transmon and the cavity, which can have very different spectral components than in the Hanh-echo experiment. A more precise understanding of the inverse-Purcell limit due to the dephasing noise would require probing the dephasing spectral density via spin-locking techniques [80].
We also estimate the cavity dephasing time limit due to the transmon thermal shot noise. In the small thermal population limit, we approximate the thermal-noise-induced cavity dephasing time as [81]
@@ Appendix B Cavity fabrication and loss characterization
To assess the decay and dephasing time scales of the bare cavity modes in the milli-Kelvin regime, we perform ringdown measurements, applying a 200 ms pulse and recording the reflected signal with a vector network analyzer. As depicted in Fig. A1(a-b), we measure the emitted power from the cavity modes in two distinct ways–phase-insensitive (PI) and phase-sensitive (PS), using the methodology introduced in [82]. The phase-insensitive trace is obtained by measuring the power amplitude (equivalent to the square sum of the cavity’s I and Q quadratures) at each time point, then averaging over many such measurements. The resultant decay trace is well described by an exponential decay function characterized by the cavity’s decay rate,
@@
Additionally, we perform a temperature sweep to investigate the temperature dependence of the internal quality factor ($Q_{0}(T)$). This allows for the extraction of the zero-temperature loss tangent ($\delta_{0}$) associated with two-level systems (TLS), the residual surface resistance ($R_{\text{res}}$), and the temperature measurement efficiency ($\beta$) using the following equation [17],
@@
The Tesla cavity geometry offers a notably low surface oxide filling factor ($F<2\times 10^{-8}$) and a large geometry factor ($G\approx 270~\Omega$), beneficial for enhancing cavity photon lifetime. At each temperature point, we perform a ring-down measurement to extract the loaded quality factor, $Q_{L}$. We then calculate the internal quality factor $Q_{0}$, from this $Q_{L}$ in conjunction with the coupling quality factor $Q_{\text{ext}}$ obtained from circle fitting at base temperature. We then fit the $Q_{0}$ vs. temperature data to Eq. B.4 and extract $F\delta_{0}$, $\beta$, and $R_{\mathrm{res}}/G$ from the fit. The factors $F$ and $G$ are obtained from finite-element method simulations using Eq. B.5 and Eq. B.6, which allows us to determine $\delta_{0}$ and $R_{\mathrm{res}}$. The relevant parameters are shown in Table A2. The measured residual surface resistance is higher than that typically observed in standard SRF cavities [78, 83, 17], suggesting the presence of residual contaminants on the cavity surface. Subsequent studies with different cavities confirmed the likely presence of such contamination in the furnace used for the cavity preparation. Based on this, elimination of these additional losses can be achieved by an extra buffered chemical polishing (BCP) etching to remove the embedded contaminants. Furthermore, we plan to reduce the TLS losses coming from surface oxide via enhanced surface treatment techniques developed previously [17, 78]. The most recent progress of SQMS (manuscript in preparation) includes demonstrating the ability to grow an air-exposure robust oxide with the drastically suppressed TLS density, which should allow a further multifold increase in the cavity quality factor. These combined improvements should allow at least an order of magnitude improvement in the energy relaxation times of both Alice and Bob modes.
@@ Appendix C Analysis of transmon-induced cavity losses
Following the methodology outlined in [84], we use finite-element simulations to calculate the loss channel participation ratios $p_{\text{bulk}}$, $p_{\text{MA}}$, $p_{\text{SA}}$, and $p_{\text{MS}}$ for the cavity modes,
@@
The accurate determination of the transmon mode’s effect on the loss channel participation ratios of the cavity mode depends critically on properly accounting for the nonlinearity-induced corrections, especially when the Lamb shift is comparable to the transmon-cavity detuning. We employ the Energy Participation Ratio (EPR) approach [85] to obtain the Lamb-shifted transmon frequency, and then renormalize the linear inductance of the Josephson junction used in the RF participation ratio simulations accordingly. This way, we are able to capture the transmon-cavity detuning precisely, allowing for a reliable estimation of the cavity’s inherited loss from the transmon and the substrate material.
@@ Appendix E Sideband-aided cavity reset operation
We realize an efficient cavity reset operation in our experiment that requires a much shorter time than via the cavity’s natural relaxation. As illustrated in Fig. A3(a), this reset operation is activated by a driven dissipation process, where the cavity population hybridizes with the transmon via the cavity-transmon sideband interaction, thus inheriting the much faster decay rate from the transmon. In our experimental regime where the sideband interaction strength is much greater than the transmon’s decay rate, the effective decay rate from $\lvert g,n\rangle$ to $\lvert e,n-1\rangle$ is around half of the transmon $\lvert f\rangle$ to $\lvert e\rangle$ decay rate [86]. Thus, the engineered driven dissipation, through the cascaded decay process of $\lvert g,n\rangle\rightarrow\lvert e,n-1\rangle\rightarrow\lvert g,n-1\rangle$, is happening at approximately the same rate as the transmon’s decay, orders of magnitude faster than the cavity’s relaxation. Using this protocol, we can reset an arbitrary initial cavity state to the vacuum state. Knowing the highest occupied Fock state $\lvert N_{\text{max}}\rangle$ in the cavity, we apply a sequence of sideband drives, going from $|g,N_{\text{max}}\rangle\leftrightarrow|f,N_{\text{max}}-1\rangle$ to $|g,1\rangle\leftrightarrow|f,0\rangle$, each lasting for a sufficiently long time comparing to the transmon’s decay time. Given its autonomous nature, this protocol does not require knowing the exact initial state, and can tolerate small detunings of the sideband drives from resonance.
@@ Appendix G Efficiency of the Sideband-Feedforward Protocol
Figure A5: Number of applied SFP operations for each Fock state. (a) Average number of applied sideband-feedforward protocol (SFP) operations during the preparation of Fock states in Alice’s mode. Ideally, the SFP is applied exactly once per sideband transition, assuming no errors occur (orange line). Experimentally, we find that the SFP is applied an average of $1.059$ times per desired sideband transition (blue line). (b) Measured average number of error syndrome states obtained during the SFP protocol for each targeted Fock state.
@@ H.1 Noise model
By assuming that the rotation angle is directly proportional to the gate time, $\mathcal{\hat{N}}_{j,k}(\theta)$ can be approximated with a second-order symmetric trotterization [87] such that
@@
where $\gamma_{T_{1}^{ge}}=1/T_{1}^{ge}$ is the transmon decay rate, and $\bar{n}_{\mathrm{th}}^{q}$ is the average thermal population. For transmon pure dephasing, we use the three-level system model presented in [88], in which pure dephasing is described by two jump operators
@@
where $\gamma_{T_{\varphi}}=1/T_{2}^{ge}-1/(2T_{1}^{ge})$ is the transmon pure dephasing rate. The factor $\sqrt{2}$ associated with the $f$ state comes as an approximation from the ratio between the noise sensitivity of the fgf-g energy difference and that of the ege-g energy difference, and more detailed discussions of the estimation of the dephasing time of the qubit can be found in [89, 90]. For the Alice-mode qudit, we write the jump operators for depolarization and pure dephasing as
@@
We also take into account the transmon readout fidelity, which introduces error to the SFP, as the correction is conditional on the ancilla transmon qubit readout. Although the confusion matrix shown in Fig. A4 is sufficient for readout error mitigation by inverting the Markov process modeling noise [91, 92], it does not distinguish between transmon-relaxation-induced readout error, which SFP can partially correct, and classification error, which SFP cannot correct. To more accurately estimate the readout-related error, we model the readout fidelity using a transmon relaxation probability, denoted as $p_{\mathrm{readout}}$, along with symmetric classification error probabilities. We use this model to simulate the experiment that measures the confusion matrix, and then optimize both the transmon relaxation probability and the classification error probabilities such that the simulated results best match the experimentally obtained confusion matrix. This yields the following results: $p_{\mathrm{readout}}(e\rightarrow g)=0.0055$, $p_{\mathrm{readout}}(f\rightarrow e)=0.0110$, and the classification error rate array
@@ Appendix I Raman-assisted beamsplitter interaction
$-\frac{E_{J}\theta_{q}^{4}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{4\Delta_{a}\Delta_{b}}\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.},$
where the second line represents the direct Alice-Bob beamsplitter interaction mediated by the 4th-order nonlinearity. By doing second-order perturbation theory to eliminate the virtual states, we find the total effective beamsplitter interaction between Alice and Bob in the $\{|10\rangle,|01\rangle\}$ subspace to be
$\hat{\mathcal{H}}_{\mathrm{VR}}^{\mathrm{BS}}/\hbar$ $\approx\left(\frac{E_{J}^{2}\theta_{q}^{8}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{2\Delta_{a}\Delta_{b}\Delta}-\frac{E_{J}\theta_{q}^{4}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{4\Delta_{a}\Delta_{b}}\right)\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.}$ (I.4)
$\approx\left(\frac{2\alpha}{\Delta}+\frac{1}{2}\right)\frac{\alpha\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{\Delta_{a}\Delta_{b}}\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.}.$ (I.5)
This clearly shows that, compared to the direct four-wave-mixing beamsplitter rate (represented by the 1/2 term), the VR scheme possesses a significant enhancement factor of $\alpha/\Delta$, making it over an order of magnitude faster than the former scheme in our experiment.
While the above derivation captures the essential interaction underlying the VR scheme, accurately modeling the system dynamics requires incorporating higher-order corrections induced by the drive or nonlinearity, which we demonstrate below. We apply a Schrieffer-Wolff transformation [93] with a generator of $\hat{\mathcal{S}}=\sum_{n=0,m=0}(\hat{\mathcal{S}}_{n,m}^{(a)}+\hat{\mathcal{S}}_{n,m}^{(b)}$), where
@@
where we define the quantities $\beta_{a,b}=\frac{E_{J}\theta_{q}^{4}\xi_{1,2}g_{a,b}}{2\Delta_{a,b}}$. From the equation above, we note that the beasmplitter rate is highly dependent on the transmon state, with detunings in the energy denominator differing by the anharmonicity for transmon $\lvert g\rangle$ and $\lvert e\rangle$ states. In general, the dispersive shifts are different for the two cavities. Thus, the only resonant term is
@@
When the transmon is in the ground state, the expression above can be approximated by the following operator equation in the regime where $\chi_{f}^{(a,b)}\ll\Delta$:
@@
The beamsplitter rate is modified because the drive detuning to the virtual states (VRBS detuning) for higher-photon-number sectors are altered by their dispersive shifts. This nonlinear correction to the beamsplitter gate has applications in universal qudit control, as discussed in Appendix N.
@@
The small but finite difference in Alice’s and Bob’s drive-induced Stark shifts results in a relative phase rotation between the Alice and Bob modes, accumulated over idling between concatenated VRBS pulses or ramping during the pulses [7]. For the error-detected VRBS swap operation, the transmon measurement after each swap can also contribute to a differential phase rotation of Alice and Bob due to their different cross-Kerr values with the readout mode (see Table A1). Without proper calibration and compensation, these accumulated phases will translate to a coherent error that reduces the fidelity of the $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$ state swap operation. In our experiment, we measure the phase rotation accumulated over a number of VRBS swap operations, by initializing the system in the $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$ state, and varying the phase $\phi$ of the final VRBS beamsplitter operation that maps the entangled state back to the single-photon Fock state in Alice. The accumulated phase is identified by the angle that maximizes the target state fidelity. We observe a linear increase of the accumulated phase with respect to the number of swap operations, where we extract the phase shift per gate to be $0.0046$ rad, see Fig. A8(b). This phase $\phi$ is then applied to the beamsplitter operation $U_{\text{BS}}(\phi)$ for the error-detected swap operation of the state $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$.
@@
Universal control of a two-qudit system requires the ability to perform arbitrary single-qudit rotations along with at least one entangling operation [94]. In our multimode architecture, single-qudit rotations $R_{d}(\vec{\theta})$ can be implemented using either sideband gates [45] or echoed conditional displacement gates [43]. The VRBS interaction discussed in the main text provides a mechanism for generating entanglement between higher Fock states as well. This operation takes a block-diagonal form, where each basis state $|m,n\rangle$ is rotated within a photon-number-conserving subspace—i.e., among states $|m^{\prime},n^{\prime}\rangle$ satisfying $m^{\prime}+n^{\prime}=m+n$. Together with single-qudit rotations, this operation enable universal control within a finite-dimensional Hilbert space constructed from $d$-dimensional qudits. For example, universal operations in the two-qudit Hilbert space can be performed by:
@@
With the results of Fig. A12, it is possible to investigate the properties of VRBS as a two-qutrit entangling gate, as a concrete example. For qutrits, it has a Schmidt rank of 9. The entangling power, $e_{p}(U)$ can be defined as the average of the linear entropy over product states $|\Psi\rangle=|\psi\rangle_{1}\otimes|\psi_{2}\rangle$ following [95]:
@@
where $\rho_{1}=\text{Tr}_{2}(U|\Psi\rangle\langle\Psi|U^{\dagger})$ is the reduced density matrix, and $U$ is the entangling gate being investigated. We estimate $e_{p}(\rm VRBS)=0.379(2)$ by sampling $10^{5}$ $|\Psi\rangle$, similar to the standard qutrit CSUM gate (${\rm CSUM_{3}}|m,n\rangle=|m,m+n\ {\rm mod}\ 3\rangle$) which is known analytically to have $e_{p}(\text{CSUM}_{3})=3/8$ [95]. This implies that VRBS is an efficient qutrit entangling gate. Further, by numerical optimization of the $SU(3)$ angles in Eq. (N.1), approximate synthesis of the CSUM3 gate from $2\leq N\leq 6$ blocks were identified with fidelities

### v3

@@ V Conclusion
The elliptical-shaped 2-cell cavity architecture demonstrated in this work can be naturally scaled by increasing the number of cells to realize larger multimode systems. These can be configured into cascaded random access quantum memories [li2025cascaded] with order-of-magnitude improvements in coherence times, where each two-cell module serves as a cache memory. Improved surface treatments [Romanenko2020, Posen2020] can further improve quality factors by an order of magnitude. Additionally, these two-cell devices can function as modular building blocks for a scalable quantum module, leveraging quantum-state router [zhou2022modular] and state-of-the-art tunable coupling devices [LuSchoelkopf2023, PRXQuantum.4.020355, Maiti2025] for the exchange of quantum information and the distribution of entanglement. This setup can be further scaled by employing remote interconnection between the scalable modules, and even across different cryogenic systems, facilitating the development of a distributed quantum network with improved coherence and transformative computational power.
@@ Appendix A System Hamiltonian parameters
Values of the system parameters are shown in Table A1. The inverse-Purcell limit on the cavity mode lifetime is determined by both the relaxation as well as the dephasing channel of the transmon, which can be approximated as [Rosenblum2023mushroom]
@@
From this, we estimate decay times of $T^{\mathrm{Purcell}}_{1a}=1.17\,$s and $T^{\mathrm{Purcell}}_{1b}=0.34\,$s, far exceeding the bare Alice and Bob lifetime limited by the cavity’s intrinsic material loss. It’s worth noting that this approximation assumes the same dephasing rate measured from the Hahn-echo experiment for the actual dressed-dephasing-induced cavity decay process. In reality, such a decay process is sensitive to dephasing noises at the detuning between the transmon and the cavity, which can have very different spectral components than in the Hanh-echo experiment. A more precise understanding of the inverse-Purcell limit due to the dephasing noise would require probing the dephasing spectral density via spin-locking techniques [Yan2013].
We also estimate the cavity dephasing time limit due to the transmon thermal shot noise. In the small thermal population limit, we approximate the thermal-noise-induced cavity dephasing time as [Clerk2007]
@@ Appendix B Cavity fabrication and loss characterization
To assess the decay and dephasing time scales of the bare cavity modes in the milli-Kelvin regime, we perform ringdown measurements, applying a 200 ms pulse and recording the reflected signal with a vector network analyzer. As depicted in Fig. A1(a-b), we measure the emitted power from the cavity modes in two distinct ways–phase-insensitive (PI) and phase-sensitive (PS), using the methodology introduced in [Read2023]. The phase-insensitive trace is obtained by measuring the power amplitude (equivalent to the square sum of the cavity’s I and Q quadratures) at each time point, then averaging over many such measurements. The resultant decay trace is well described by an exponential decay function characterized by the cavity’s decay rate,
@@
Additionally, we perform a temperature sweep to investigate the temperature dependence of the internal quality factor ($Q_{0}(T)$). This allows for the extraction of the zero-temperature loss tangent ($\delta_{0}$) associated with two-level systems (TLS), the residual surface resistance ($R_{\text{res}}$), and the temperature measurement efficiency ($\beta$) using the following equation [Romanenko2020],
@@
The Tesla cavity geometry offers a notably low surface oxide filling factor ($F<2\times 10^{-8}$) and a large geometry factor ($G\approx 270~\Omega$), beneficial for enhancing cavity photon lifetime. At each temperature point, we perform a ring-down measurement to extract the loaded quality factor, $Q_{L}$. We then calculate the internal quality factor $Q_{0}$, from this $Q_{L}$ in conjunction with the coupling quality factor $Q_{\text{ext}}$ obtained from circle fitting at base temperature. We then fit the $Q_{0}$ vs. temperature data to Eq. B.4 and extract $F\delta_{0}$, $\beta$, and $R_{\mathrm{res}}/G$ from the fit. The factors $F$ and $G$ are obtained from finite-element method simulations using Eq. B.5 and Eq. B.6, which allows us to determine $\delta_{0}$ and $R_{\mathrm{res}}$. The relevant parameters are shown in Table A2. The measured residual surface resistance is higher than that typically observed in standard SRF cavities [Posen2020, Romanenko2017, Romanenko2020], suggesting the presence of residual contaminants on the cavity surface. Subsequent studies with different cavities confirmed the likely presence of such contamination in the furnace used for the cavity preparation. Based on this, elimination of these additional losses can be achieved by an extra buffered chemical polishing (BCP) etching to remove the embedded contaminants. Furthermore, we plan to reduce the TLS losses coming from surface oxide via enhanced surface treatment techniques developed previously [Romanenko2020, Posen2020]. The most recent progress of SQMS (manuscript in preparation) includes demonstrating the ability to grow an air-exposure robust oxide with the drastically suppressed TLS density, which should allow a further multifold increase in the cavity quality factor. These combined improvements should allow at least an order of magnitude improvement in the energy relaxation times of both Alice and Bob modes.
@@ Appendix C Analysis of transmon-induced cavity losses
Following the methodology outlined in [Ganjam2024], we use finite-element simulations to calculate the loss channel participation ratios $p_{\text{bulk}}$, $p_{\text{MA}}$, $p_{\text{SA}}$, and $p_{\text{MS}}$ for the cavity modes,
@@
The accurate determination of the transmon mode’s effect on the loss channel participation ratios of the cavity mode depends critically on properly accounting for the nonlinearity-induced corrections, especially when the Lamb shift is comparable to the transmon-cavity detuning. We employ the Energy Participation Ratio (EPR) approach [Minev2021] to obtain the Lamb-shifted transmon frequency, and then renormalize the linear inductance of the Josephson junction used in the RF participation ratio simulations accordingly. This way, we are able to capture the transmon-cavity detuning precisely, allowing for a reliable estimation of the cavity’s inherited loss from the transmon and the substrate material.
@@ Appendix E Sideband-aided cavity reset operation
We realize an efficient cavity reset operation in our experiment that requires a much shorter time than via the cavity’s natural relaxation. As illustrated in Fig. A3(a), this reset operation is activated by a driven dissipation process, where the cavity population hybridizes with the transmon via the cavity-transmon sideband interaction, thus inheriting the much faster decay rate from the transmon. In our experimental regime where the sideband interaction strength is much greater than the transmon’s decay rate, the effective decay rate from $\lvert g,n\rangle$ to $\lvert e,n-1\rangle$ is around half of the transmon $\lvert f\rangle$ to $\lvert e\rangle$ decay rate [Lu2017]. Thus, the engineered driven dissipation, through the cascaded decay process of $\lvert g,n\rangle\rightarrow\lvert e,n-1\rangle\rightarrow\lvert g,n-1\rangle$, is happening at approximately the same rate as the transmon’s decay, orders of magnitude faster than the cavity’s relaxation. Using this protocol, we can reset an arbitrary initial cavity state to the vacuum state. Knowing the highest occupied Fock state $\lvert N_{\text{max}}\rangle$ in the cavity, we apply a sequence of sideband drives, going from $|g,N_{\text{max}}\rangle\leftrightarrow|f,N_{\text{max}}-1\rangle$ to $|g,1\rangle\leftrightarrow|f,0\rangle$, each lasting for a sufficiently long time comparing to the transmon’s decay time. Given its autonomous nature, this protocol does not require knowing the exact initial state, and can tolerate small detunings of the sideband drives from resonance.
@@ Appendix G Efficiency of the Sideband-Feedforward Protocol
Figure A5: Number of applied SFP operations for each Fock state. (a) Average number of applied sideband-feedforward protocol (SFP) operations during the preparation of Fock states in Alice’s mode. Ideally, the SFP is applied exactly once per sideband transition, assuming no errors occur (orange line). Experimentally, we find that the SFP is applied an average of $1.059$ times per desired sideband transition (blue line). (b) Measured average number of error syndrome states obtained during the SFP protocol for each targeted Fock state. (c) Measured success rate of the parity filter following the SFP operation. The filter rejects states with the incorrect parity, and the observed success rate is consistent with the Fock state preparation fidelity achieved using the SFP protocol alone.
@@ H.1 Noise model
By assuming that the rotation angle is directly proportional to the gate time, $\mathcal{\hat{N}}_{j,k}(\theta)$ can be approximated with a second-order symmetric trotterization [Suzuki1991] such that
@@
where $\gamma_{T_{1}^{ge}}=1/T_{1}^{ge}$ is the transmon decay rate, and $\bar{n}_{\mathrm{th}}^{q}$ is the average thermal population. For transmon pure dephasing, we use the three-level system model presented in [Li2012], in which pure dephasing is described by two jump operators
@@
where $\gamma_{T_{\varphi}}=1/T_{2}^{ge}-1/(2T_{1}^{ge})$ is the transmon pure dephasing rate. The factor $\sqrt{2}$ associated with the $f$ state comes as an approximation from the ratio between the noise sensitivity of the fgf-g energy difference and that of the ege-g energy difference, and more detailed discussions of the estimation of the dephasing time of the qubit can be found in [Koch2007, Groszkowski2018]. For the Alice-mode qudit, we write the jump operators for depolarization and pure dephasing as
@@
We also take into account the transmon readout fidelity, which introduces error to the SFP, as the correction is conditional on the ancilla transmon qubit readout. Although the confusion matrix shown in Fig. A4 is sufficient for readout error mitigation by inverting the Markov process modeling noise [Geller2020, Peters2023], it does not distinguish between transmon-relaxation-induced readout error, which SFP can partially correct, and classification error, which SFP cannot correct. To more accurately estimate the readout-related error, we model the readout fidelity using a transmon relaxation probability, denoted as $p_{\mathrm{readout}}$, along with symmetric classification error probabilities. We use this model to simulate the experiment that measures the confusion matrix, and then optimize both the transmon relaxation probability and the classification error probabilities such that the simulated results best match the experimentally obtained confusion matrix. This yields the following results: $p_{\mathrm{readout}}(e\rightarrow g)=0.0055$, $p_{\mathrm{readout}}(f\rightarrow e)=0.0110$, and the classification error rate array
@@ Appendix I Raman-assisted beamsplitter interaction
$-\frac{E_{J}\theta_{q}^{4}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{\Delta_{a}\Delta_{b}}\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.},$
where the third line represents the direct Alice-Bob beamsplitter interaction mediated by the 4th-order nonlinearity. By doing second-order perturbation theory to eliminate the virtual states, we find the total effective beamsplitter interaction between Alice and Bob in the $\{|10\rangle,|01\rangle\}$ subspace to be
$\hat{\mathcal{H}}_{\mathrm{VR}}^{\mathrm{BS}}/\hbar$ $\approx\left(\frac{E_{J}^{2}\theta_{q}^{8}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{2\Delta_{a}\Delta_{b}\Delta}-\frac{E_{J}\theta_{q}^{4}\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{\Delta_{a}\Delta_{b}}\right)\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.}$ (I.4)
$\approx 2\left(\frac{\alpha}{\Delta}+1\right)\frac{\alpha\xi_{1}\xi^{*}_{2}g_{a}g_{b}}{\Delta_{a}\Delta_{b}}\hat{a}\hat{b}^{\dagger}+\mathrm{h.c.}.$ (I.5)
This clearly shows that, compared to the direct four-wave-mixing beamsplitter rate (represented by the 1 term), the VR scheme possesses a significant enhancement factor of $\alpha/\Delta$, making it over an order of magnitude faster than the former scheme in our experiment.
While the above derivation captures the essential interaction underlying the VR scheme, accurately modeling the system dynamics requires incorporating higher-order corrections induced by the drive or nonlinearity, which we demonstrate below. We apply a Schrieffer-Wolff transformation [Shavitt1980] with a generator of $\hat{\mathcal{S}}=\sum_{n=0,m=0}(\hat{\mathcal{S}}_{n,m}^{(a)}+\hat{\mathcal{S}}_{n,m}^{(b)}$), where
@@
where we define the quantities $\beta_{a,b}=\frac{E_{J}\theta_{q}^{4}\xi_{1,2}g_{a,b}}{2\Delta_{a,b}}$. From the equation above, we note that the beasmplitter rate is highly dependent on the transmon state, with detunings in the energy denominator differing by the anharmonicity for transmon $\lvert g\rangle$ and $\lvert e\rangle$ states. In general, the dispersive shifts are different for the two cavities. When the transmon is in the ground state, the resonant interaction is
@@
This expression can be further simplied in the regime where $\chi_{f}^{(a,b)}\ll\Delta$:
@@
Intuitively, the beamsplitter rate is dependent on the cavity photon number because the drive detuning to the virtual states (VRBS detuning) for higher-photon-number sectors is altered by their dispersive shifts. This nonlinear correction to the beamsplitter gate has applications in universal qudit control, as discussed in Appendix N.
@@
The small but finite difference in Alice’s and Bob’s drive-induced Stark shifts results in a relative phase rotation between the Alice and Bob modes, accumulated over idling between concatenated VRBS pulses or ramping during the pulses [LuSchoelkopf2023]. For the error-detected VRBS swap operation, the transmon measurement after each swap can also contribute to a differential phase rotation of Alice and Bob due to their different cross-Kerr values with the readout mode (see Table A1). Without proper calibration and compensation, these accumulated phases will translate to a coherent error that reduces the fidelity of the $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$ state swap operation. In our experiment, we measure the phase rotation accumulated over a number of VRBS swap operations, by initializing the system in the $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$ state, and varying the phase $\phi$ of the final VRBS beamsplitter operation that maps the entangled state back to the single-photon Fock state in Alice. The accumulated phase is identified by the angle that maximizes the target state fidelity. We observe a linear increase of the accumulated phase with respect to the number of swap operations, where we extract the phase shift per gate to be $0.0046$ rad, see Fig. A8(b). This phase $\phi$ is then applied to the beamsplitter operation $U_{\text{BS}}(\phi)$ for the error-detected swap operation of the state $(\lvert 01\rangle+\lvert 10\rangle)/\sqrt{2}$.
@@
Universal control of a two-qudit system requires the ability to perform arbitrary single-qudit rotations along with at least one entangling operation [Braunstein2005CV]. In our multimode architecture, single-qudit rotations $R_{d}(\vec{\theta})$ can be implemented using either sideband gates [Liu2021qudit] or echoed conditional displacement gates [You2024crosstalk]. The VRBS interaction discussed in the main text provides a mechanism for generating entanglement between higher Fock states as well. This operation takes a block-diagonal form, where each basis state $|m,n\rangle$ is rotated within a photon-number-conserving subspace—i.e., among states $|m^{\prime},n^{\prime}\rangle$ satisfying $m^{\prime}+n^{\prime}=m+n$. Together with single-qudit rotations, this operation enable universal control within a finite-dimensional Hilbert space constructed from $d$-dimensional qudits. For example, universal operations in the two-qudit Hilbert space can be performed by:
@@
With the results of Fig. A12, it is possible to investigate the properties of VRBS as a two-qutrit entangling gate, as a concrete example. For qutrits, it has a Schmidt rank of 9. The entangling power, $e_{p}(U)$ can be defined as the average of the linear entropy over product states $|\Psi\rangle=|\psi\rangle_{1}\otimes|\psi_{2}\rangle$ following [wang2003entangling]:
@@
where $\rho_{1}=\text{Tr}_{2}(U|\Psi\rangle\langle\Psi|U^{\dagger})$ is the reduced density matrix, and $U$ is the entangling gate being investigated. We estimate $e_{p}(\rm VRBS)=0.379(2)$ by sampling $10^{5}$ $|\Psi\rangle$, similar to the standard qutrit CSUM gate (${\rm CSUM_{3}}|m,n\rangle=|m,m+n\ {\rm mod}\ 3\rangle$) which is known analytically to have $e_{p}(\text{CSUM}_{3})=3/8$ [wang2003entangling]. This implies that VRBS is an efficient qutrit entangling gate. Further, by numerical optimization of the $SU(3)$ angles in Eq. (N.1), approximate synthesis of the CSUM3 gate from $2\leq N\leq 6$ blocks were identified with fidelities